Gamma rays are photons with very high energy. How many visible-light photons with a wavelength of 500 nm would you need to equal the energy of a gamma-ray photon with energy 4.1x10^-13J

A) 1.0 × 106

B) 1.4 × 108

C) 6.2 × 109

D) 3.9 × 103
Answer #1

Energy of a photon = hc/λ

where h =6.62606957 × 10-34

C= speed of light = 3*10^8 m/s

and wavelength λ = 500 nm

energy of 1 photon =(6.62606957*10^-34 *3*10^8)/(500*10-9)

=3.956*10^-19 J

No of photons needed =4.1*10-13/(3.956*10^-19)=1.03*10^6

ans is A